OpenStudy (anonymous):

Estimate the square root to the nearest integer Sqaure root of 25 A. 6 B. 5 C. 7 D. 4

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (freckles):

25 is a perfect square

OpenStudy (freckles):

estimation not really needed

OpenStudy (anonymous):

So 5

OpenStudy (freckles):

yes 5^2=5(5)=25

OpenStudy (anonymous):

Thx

OpenStudy (freckles):

so sqrt(25)=sqrt(5*5)=5